OpenStudy (anonymous):

evaluate the given integral 5.4e^(-4x+5)dx

OpenStudy (blurbendy):

factor out 5.4, so you have 5.4 (integral) e^(-4x+5)dx, then use substitution

OpenStudy (anonymous):

So I know that u would be -4x+5 and du would be -4 dx but I dont know what to do after that

OpenStudy (blurbendy):

dx = -1/4du -1/4 * 5.4 (integral) e^u take the integral of e^u

OpenStudy (dumbcow):

right so then plug in for u and du dx = du/-4 \[-\frac{5.4}{4} \int\limits e^{u} du\]

OpenStudy (anonymous):

Ok thanks!

OpenStudy (blurbendy):

np